3. Derive homogenous wave equation for Hcomponent of an EM wave in a lossless, linear and isotropic medium.

Electromagnetic
Answer #1

ĒXĦ => P S xH given the propagation of p] & 1 => 1 A 2 - Gre Inadoni El Teruntunited wamo LOCIVO - RET medium 1 12 Thoden ic] ET When HEL = HE 2. ] modior modium 2 1 (HR, Hp Opposito dire divociion Sobradi boin) Hi - He 2 HT [E/H = 2 H = Ein ChangeTOP SONUN It Ei = Er wo get some direction then it satisfied XEiL 22.>Ni Er=ſeet POSTive IZ go Hij HR = TH Hi HR = TE Hi ERXa at EM, WavÓ In non conducting med en Dod=ev V.B=0 Dxe -2B VxH = 5+2D at electric *Assuming fields que in free spaco C=o), aӘНу . 22 Hy DPEx = I D2E> d22 V2 at 2+2 222 L EM wave put EX2 50 t), HY (Z3Lt) It wo best any time varying field inpes any reӘНу . 22 Hy DPEx = I D2E> d22 V2 at 2+2 222 L EM wave put EX2 50 t), HY (Z3Lt) It wo best any time varying field inpes any re. Hy = P_ Ext6-182 - B EX CBZ Hy = -1 Joy 1-jp Ext eiBz +jB E01B2] WH WH Hyt ну Hy = Hy tow]? + Hyºe JB2 Where Hy + = B. ExtJ Ex(2) =n Exto JB2 - Ext 6-1BZ Ну (z) #4821 n - En pa an with complex representation wows/m2 Pavo = = Re EXH* Exfasz x Ext:. Hy = P_ Ext6-182 - B EX CBZ Hy = -1 Joy 1-jp Ext eiBz +jB E01B2] WH WH Hyt ну Hy = Hy tow]? + Hyºe JB2 Where Hy + = B. ExtJ Ex(2) =n Exto JB2 - Ext 6-1BZ Ну (z) #4821 n - En pa an with complex representation wows/m2 Pavo = = Re EXH* Exfasz x Ext: